# Performance Assessment of Universal Machine Learning Interatomic Potentials: Challenges and Directions for Materials’ Surfaces

### v1

@@ II.4.1 Universal vs NequIP/MTP
Surface total energy
[…]
Surface energy
[…]
Type
MACE 0.0471 0.0488 0.3974 UIP
M3GNet 0.0227 0.0979 0.6920 UIP
CHGNet 0.0791 0.0386 0.5128 UIP
MTP [JPCA] 0.0369 0.0480 0.2715 LRIP
MTP [$*{}^{*}$] 0.0519 0.4281 2.7932 LRIP
NequIP [JPCA] 0.0372 0.0181 0.2097 GNN
NequIP [$*{}^{*}$] 0.0316 0.0762 0.6511 GNN
MACE [FT] 0.0367 0.0496 0.2183 UIP-FT
As demonstrated in Fig. 6, for most elements where UIPs show an increased error, MTP [JPCA] and NequIP [JPCA] demonstrate high-accuracy predictions. Predictions for systems such as Mo and Si show great improvement from the specialized MLIPs. Systems containing Ge on the other hand show worse performance for bulk structures while slightly improved performance for surface structures.
II.4.2 Universal (zero-shot) vs Universal (fine-tuned)
Since UIPs seem to not be great at dealing with surfaces, we now address the question: “What can we do to improve them?”. One route to improving general models such as the UIPs is fine-tuning, which consists of continuing the training of these models however with a different dataset, tailored to make these universal models more specialized in particular tasks. In our case, we can use the training and test sets of ref. [25] to fine-tune these UIPs. By introducing a more diverse set of structures for each chemical element, we expect that the UIP will improve in predicting surface (total) energies while taking advantage of what it has already learned from the MPtrj dataset. We choose to fine-tune the MACE model, given the availability and ease of use of the corresponding implementation. We fine-tune our MACE model for another 200 epochs using the training set from ref. [25]. The fine-tuned model is called MACE [FT], as shown in the last line of Fig. 6 and Table 1.
After fine-tuning, we observe that the prediction of bulk total energies largely improved for most systems. Is noticeable how fine-tuning should be performed with care. As discussed before, both MTP and NequIP did not show great performance over bulk structures composed of germanium (Ge). After fine-tuning, MACE [FT] decreased its performance on germanium (Ge) bulk structures. The performance on surface (total) energies improved for most elements. On nickel (Ni) the RMSE slightly increased, and for copper (Cu) the increase was more significant. Overall, the fine-tuned universal model represents an important improvement: for the selected systems, the fine-tuned MACE model is of similar performance to specialized MLIPs.
@@ III Conclusions
We have assessed one of the main questions for working with MLIPs: given a new problem, what is the best and most efficient way to tackle it? The answer follows at least two steps, evaluating if MLIPs are accurate enough and if not, if it is better to train a specialized MLIP from scratch or to fine-tune a universal foundational model. From our results for surface energies, we see that the total energies for surface geometries are modestly accurate, however, not good enough for specific properties. Therefore, as expected, the most important aspect is the domain distance of the training dataset to the target space, composed of both chemistry and structure.
In terms of efficiency, fine-tuning UIPs can greatly accelerate the training by incorporating alchemical transferable knowledge from the large bulk dataset (in this case, MPtrj), thus requiring only a modest dataset to achieve sufficient accuracy in specialized tasks. The combination of active learning strategies using cheap UIP trajectories to generate synthetic datasets for pre-training MLIPs can also greatly increase the efficiency of generating new datasets and expanding current ones [29, 30].
The performance of MLIPs for surfaces is important because, in materials science, a plethora of different phenomena take place on the surface of materials. Examples are adsorption and absorption of molecules and gases, catalysis, electrochemical processes such as water splitting, deposition and growth of new materials and phases such as heterostructures and nanoparticles. The generalization including both extensivity and transferability of MLIPs allows for simulations and understanding of ever larger and complex systems at an efficient computational cost. This enables the study of nanoscale and quantum materials systems such as 2D materials and their twisted interfaces, as well as more general global challenges involving molecules and materials, such as in energy, health, and the environment [31, 32, 33, 34, 35, 36, 37, 38, 39, 40]. As these models become more used, understanding their application limits and broadening them is of great importance to accelerate materials science research and applications of these models to different systems and use cases.
@@ IV.1 Datasets
The Materials Project surface dataset [24] was accessed on Feb. 2023 through Materials Project’s API [43]. We call $*{}^{*}$ the MP data that is filtered for the following set of elements from the MPtrj [44] dataset: Ni (96 structures), Cu (55 structures), Li (83 structures), Mo (25 structures), Si (559 structures), Ge (225 structures). Accordingly, for MTP and NequIP, a different model is fit for each element. The entire $*{}^{*}$ dataset is used for training and the test is performed on the MP surfaces dataset. The so-called JPCA dataset is the one made available from ref. [25].
IV.2 Training specialized potentials
IV.2.1 NequIP models
[…]
IV.2.2 MTP models
The moment tensor potential (MTP) models construct a contracted rotationally invariant representation of the atomic local environments with tensors from a set of basis functions, these are used to build a linear expansion of the potential energy as a function of the atomic representation [27]. The MTP models trained with the $*{}^{*}$ dataset are trained using the same radius cutoff as the models in ref. [25] for each element, while other hyperparameters are kept as the default from the maml package [46], while the so-called MTP [JPCA] models are the models made available from ref. [25]. Trained MTP models are available as Supporting Information.
IV.3 Fine-tuning MACE model
@@
We fine-tune the MACE model with the aid of the training scripts provided with the MACE code package. The initial model for fine-tuning MACE is the MACE-MP-0 model trained on ref. [19], used throughout the text as MACE. The data used for fine-tuning is available from ref. [25]. Further details and command options, as well as the trained model, are made available as Supporting Information.
@@
Dataset exploration: surface energy data distribution; Evaluation of UIPs: correlation between UIPs’ predictions, understanding the errors by system chemistry; Comparison between UIPs, specialized MLIPs, and fine-tuned UIP: parity plot of all models. All trained models are available as files with instructions on their use in a compressed .zip file.

### v2

@@ II.4.3 Surface relative energy order
In addition to the surface properties, we performed molecular dynamics (MD) runs for the higher RMSE surfaces of each element. To assess the stability of the trajectories from the MLIPs considered in this work, we carried out NVT MD simulations at $T=0.8\times\text{melting point}$ for the higher RMSE surfaces of each corresponding element for $30\;\rm ps$. For all UIPs and MACE [FT], the dynamics were stable with no qualitative difference between the MLIPs.
@@ III Conclusions
We have assessed one of the main questions for working with MLIPs: given a new problem, what is the best and most efficient way to tackle it? The answer follows at least two steps, evaluating if MLIPs are accurate enough and if not, if it is better to train a specialized MLIP from scratch or to fine-tune a universal foundational model. From our results for surface energies, we see that the total energies for surface geometries are modestly accurate, however, not good enough for specific properties. Therefore, as expected, the most important aspect is the coverage of the training dataset to the target space, composed of both chemistry and structure. The investigation of what are the most efficient strategies to cover this space is an important question for future studies, as UIPs can display low errors in structures close to equilibrium [33], suggesting a connection between out-of-domain distance and prediction errors [11, 29].
In terms of efficiency, fine-tuning UIPs can greatly accelerate the training by incorporating alchemical transferable knowledge from the large bulk dataset (in this case, MPtrj), thus requiring only a modest dataset to achieve sufficient accuracy in specialized tasks. The combination of active learning strategies using cheap UIP trajectories to generate synthetic datasets for pre-training MLIPs can also greatly increase the efficiency of generating new datasets and expanding current ones [27, 34]. Making use of already existing datasets, although at different fidelity levels, can be explored in multi-fidelity or multi-modality approaches, recently shown to be beneficial and accelerating fine-tuning more than 100$\times$ [35]. Additionally, there is still plenty of exploration of potential gains related to the architectural design of the models. Beyond graph Atomic Cluster Expansions [15], models based on transformers such as EquiformerV2 [36] display improved data efficiency, and MatterSim models [24] using this architecture are also comparatively more accurate.
The performance of MLIPs for surfaces is important because, in materials science, a plethora of different phenomena take place on the surface of materials. Examples are adsorption and absorption of molecules and gases, catalysis, electrochemical processes such as water splitting, deposition and growth of new materials and phases such as heterostructures and nanoparticles. The generalization including both extensivity and transferability of MLIPs allows for simulations and understanding of ever larger and complex systems at an efficient computational cost. This enables the study of nanoscale and quantum materials systems such as 2D materials and their twisted interfaces, as well as more general global challenges involving molecules and materials, such as in energy, health, and the environment [37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48]. As these models become more used, understanding their application limits and broadening them is of great importance to accelerate materials science research and applications of these models to different systems and use cases.
@@ IV.1 Datasets
The Materials Project surface dataset (crystalium) [26] was accessed on Feb. 2023 through Materials Project’s API [51].
@@
We call MPtrj the MP data that is filtered for unary structure from the following set of elements from the MPtrj [52] dataset: Ni (96 structures), Cu (55 structures), Li (83 structures), Mo (25 structures), Si (559 structures), Ge (225 structures). Accordingly, for MTP and NequIP, a different model is fit for each element. The entire MPtrj dataset is used for training and the test is performed on the MP surfaces dataset. Further details can be found in the Supporting Information.
The so-called [Ong21] dataset is made available from ref. [28] and contains structures for Ni (263/31 structures), Cu (262/31 structures), Li (241/29 structures), Mo (194/23 structures), Si (214/25 structures), Ge (228/25 structures) for training/testing tasks. Further details can be found on ref. [28] and the maml package on https://github.com/materialsvirtuallab/maml.
IV.2 Universal potentials
IV.2.1 Model’s versions
Table 2 shows the details and versions of the UIPs tested in this work.
Table 2: Details and version of UIPs tested.
Model Version Model size Dataset Data size
M3GNet [18] 2021.2.8-DIRECT [27] 1.1M MPF [18] 185.6k
CHGNet [19] v0.3.0 412.5k MPtrj [52] 1.58M
MACE [20] 2024.01.07-128-L2 5.7M MPtrj [52] 1.58M
IV.2.2 Fine-tuning
@@
We fine-tune the MACE model with the aid of the training scripts provided with the mace code package available from https://github.com/ACEsuit/mace. The initial model for fine-tuning MACE is the MACE-MP-0 (for $L=2$) model, as specified in Table 2, used throughout the text as MACE, available from https://huggingface.co/cyrusyc/mace-universal (2024-01-07-mace-128-L2_epoch-199.model). The data used for fine-tuning is available from ref. [28]. Fine-tuning was performed for a maximum of 200 epochs, training was stopped if the validation metric did not improve for 50 epochs. Further details, as well as the trained model, are made available as Supporting Information.
Fine-tuning the M3GNet model is performed with aid of the matgl code package available from https://github.com/materialsvirtuallab/matgl. The initial model for fine-tuning M3GNET is the M3GNET-DIRECT model trained in ref. [27], as specified in Table 2, and used throughout the text as M3GNET. The data used for fine-tuning is available from ref. [28]. Fine-tuning was performed for a maximum of 200 epochs, training was stopped if the validation metric did not improve for 50 epochs. Further details, as well as the trained model, are made available as Supporting Information.
Fine-tuning for both MACE and M3GNet was performed using a single NVIDIA Tesla P100 16 GB GPU, taking 14 hours on average for MACE and 3 hours on average for M3GNet using 1402 structures in both cases.
IV.3 Specialized potentials
IV.3.1 MACE from scratch
We trained a MACE model using the training scripts provided with the mace code package available from https://github.com/ACEsuit/mace. Hyperparameters for training MACE were set to spherical expansion of up to $l_{\rm max}=3$, and 4-body messages in each layer; 128-channel dimension for tensor decomposition; radius cutoff of 6 Å; 10 Bessel functions for radial basis; three hidden layers of 64 hidden units; and $L=2$ resulting in the 128x0e+128x1o+128x2e irreducible representations. The weights for the Huber loss considering energy, forces, and stress were set to $\lambda_{E}=1$, $\lambda_{F}=10$, and $\lambda_{\sigma}=100$, following ref. [22]. All other hyperparameters were the same as in ref. [22]. The data used for training is available from ref. [28]. The trained models are available as Supporting Information.
IV.3.2 M3GNet from scratch
We trained a M3GNet model using the training scripts provided with the matgl code package available from https://github.com/materialsvirtuallab/matgl. Hyperparameters for training M3GNet were set to a 1:1:0.1 weight ratio for energy (eV atom-1), force (eV Å-1), and stress (GPa) in a Huber loss function with $\delta=0.01$, an Adam optimizer with initial rate of $10^{-3}$ and a cosine decay to 1% of the original value in 100 epochs. The training was performed for 200 epochs with 50 epochs of patience over the validation total loss. The number of hidden units, dimension of node, and edge embeddings were set to 128. The radius cutoff was set to $5$, and all other hyperparameters were set to be the same as in M3GNet trained in ref. [27]The data used for training is available from ref. [28]. The trained models are available as Supporting Information.
IV.3.3 MTP models
The moment tensor potential (MTP) models construct a contracted rotationally invariant representation of the atomic local environments with tensors from a set of basis functions, these are used to build a linear expansion of the potential energy as a function of the atomic representation [31]. The so-called MTP [Ong21] models are made available from ref. [28] through the maml package on https://github.com/materialsvirtuallab/maml. The MTP models trained with the MPtrj dataset are trained using the set of hyperparameters selected via grid search. Further details and the trained models are available as Supporting Information.
IV.3.4 NequIP models
Neural Equivariant Interatomic Potential (NequIP) is a highly accurate approach to generate MLIP models with outstanding data efficiency. The architecture of the model yields an E(3)-equivariant graph neural network in which the features are updated in a message passing scheme and the rotation equivariance is achieved by convolution filters via tensor product of radial functions and spherical harmonics [16, 53]. In this case, the convolution operation compromises the atoms close to a central atom within a cutoff radius which results in semi-local environments. Herein we construct NequIP models from MPtrj and [Ong21] datasets. The model is trained using the nequip package available from https://github.com/mir-group/nequip. The set of hyperparameters is selected via grid search. Further details and the trained models are available as Supporting Information.
@@ IV.3.4 NequIP models
Dataset exploration: surface energy data distribution; Evaluation of UIPs: correlation between UIPs’ predictions, understanding the errors by system chemistry; Comparison between UIPs, specialized MLIPs, and fine-tuned UIP: parity plot and root mean squared errors table of all models; Performance and computational cost concerning fine-tuning training volume; Hyperparameter optimization of specialized IPs. All trained models and code are available as files with instructions on their use in the following repository [54]: https://doi.org/10.5281/zenodo.11391989.